solve the following differential equations -


sin-1(dy/dx)= x+y


(1+y)2dx=(tan-1y-x)dy


(D-1)2(D2+1)2 Y= sin2 (x/2) + ex + x


2x2y(d2y/dx2) + 4y2= x2 ( dy/dx)2 +2xy (dy/dx)

1
Expert's answer
2022-01-13T07:32:56-0500

1)

"y'=sin(x+y)"

"x+y=v"

"v'-1=y'"

"v'-1=sinv"

"\frac{dv}{1+sinv}=dx"


"\frac{dv}{1+sinv}=\int \frac{sec^2(v/2)}{(tan(v/2)+1)^2}dv=|u=tan(v/2)+1,dv=\frac{2}{sec^2(v/2)}du|="


"=\int 2 du/u^2=-2/u=-\frac{2}{tan(v/2)+1}"


"-\frac{2}{tan((x+y)/2)+1}=x+c"



2)

"R=(y+1)^2,S=x-tan^{-1}y"

find integrating factor "\mu(y)" such that

"\mu R+y'\mu S=0" is exact

then:

"\frac{\partial}{\partial y}(\mu R)=\frac{\partial}{\partial x}(\mu S)"


"d\mu /dy(y+1)^2+2(y+1)\mu =\mu"


"\frac{d\mu /dy}{\mu}=\frac{-2y-1}{(y+1)^2}"


"ln\mu =-2ln(y+1)-\frac{1}{y+1}"

integrating factor:

"\mu(y)=\frac{e^{-1/(y+1)}}{(y+1)^2}"


then:

"e^{-1/(y+1)}+\frac{(e^{-1/(y+1)}(x-tan^{-1}y))y'}{(y+1)^2}=0"


"P(x,y)=e^{-1/(y+1)},Q(x,y)=\frac{e^{-1/(y+1)}(x-tan^{-1}y)}{(y+1)^2}"


"f(x,y)=\int P(x,y)dx=\int e^{-1/(y+1)}dx=xe^{-1/(y+1)}+g(y)"


"\frac{\partial f(x,y)}{\partial y}=\frac{xe^{-1/(y+1)}}{(y+1)^2}+\frac{dg(y)}{dy}"


"\frac{\partial f(x,y)}{\partial y}=Q(x,y)"


"\frac{xe^{-1/(y+1)}}{(y+1)^2}+\frac{dg(y)}{dy}=\frac{e^{-1/(y+1)}(x-tan^{-1}y)}{(y+1)^2}"


"\frac{dg(y)}{dy}=-\frac{e^{-1/(y+1)}tan^{-1}y}{(y+1)^2}"


"g(y)=\int -\frac{e^{-1/(y+1)}tan^{-1}y}{(y+1)^2}dy=-e^{-1/(y+1)}-\frac{1}{2}ie^{-1/2+i/2}Ei(-\frac{1}{y+1}+\frac{1-i}{2})+"


"+\frac{1}{2}ie^{-1/2-i/2}Ei(-\frac{1}{y+1}+\frac{1+i}{2})--e^{-1/(y+1)}(tan^{-1}y-1)"


"f(x,y)=xe^{-1/(y+1)}-e^{-1/(y+1)}-\frac{1}{2}ie^{-1/2+i/2}Ei(-\frac{1}{y+1}+\frac{1-i}{2})+"


"+\frac{1}{2}ie^{-1/2-i/2}Ei(-\frac{1}{y+1}+\frac{1+i}{2})--e^{-1/(y+1)}(tan^{-1}y-1)"


solution:

"f(x,y)=c"


"xe^{-1/(y+1)}-e^{-1/(y+1)}-\frac{1}{2}ie^{-1/2+i/2}Ei(-\frac{1}{y+1}+\frac{1-i}{2})+"

"+\frac{1}{2}ie^{-1/2-i/2}Ei(-\frac{1}{y+1}+\frac{1+i}{2})--e^{-1/(y+1)}(tan^{-1}y-1)=c"


where Ei is exponential integral:

"Ei(x)=\int_{-\infin}^x e^t dt/t"


3)

characteristic equation:

"(k-1)^2(k^2+1)^2=0"

"k_{1,2}=1"

"k_{3,4,5,6}=\pm i"

"y_c=c_1e^x+c_2xe^x+c_3cosx+c_4sinx+c_5xcosx+c_6xsinx"


for particular solution:

"y_{p1}=\frac{e^x}{(1+1)^2}\frac{x^2}{2!}=x^2e^x/4"


"sin^2 (x/2)=\frac{1-cosx}{2}"


"y_{p2}=-\frac{1}{2(-2-2D)(D^2+1)}\frac{x}{2}\int cosxdx=\frac{1}{8(1+D)(D^2+1)}xsinx"


"\frac{1}{(D^2+1)}xsinx=\frac{1}{(D^2+1)}\cdot I.P.(xe^{ix})= I.P.e^{ix}\frac{1}{2Di}(1-\frac{D}{2i}+...)x="


"I.P.e^{ix}\frac{1}{2Di}(x-\frac{1}{2i})= I.P.e^{ix}\frac{1}{2i}(x^2/2-\frac{x}{2i})="


"= I.P.(x^2/2-\frac{x}{2i})(cosx+isinx)=xcosx/2+x^2sinx/2"


where I.P. is Imaginary part


"y_{p2}=\frac{1}{8(1+D)}(xcosx/2+x^2sinx/2)"


"\frac{1}{(1+D)}xcosx=R.P.e^{ix}\frac{1}{(1+D+i)}x=R.P.e^{ix}(\frac{1-i}{2}+iD/2-...)x="


"=R.P.((x-ix+i))(cosx+isinx)\/2=(xcosx+(x-1)sinx)\/2"


where R.P. is real part


"\frac{1}{(1+D)}x^2sinx=I.P.e^{ix}(\frac{1-i}{2}+iD/2-(1+i)D^2/4+...)x^2="


"=I.P.(cosx+isinx)((x^2-ix^2)\/2+ix-(1+i)\/2)="


"=-x^2cosx\/2+xcosx-cosx\/2+x^2sinx\/2-sinx\/2"


"y_{p2}=\frac{1}{32}((x-2)sinx -(x^2+1)cosx+3xcosx+x^2sinx)"


"y_{p3}=\frac{1}{(D-1)^2(D^2+1)^2}(x+1/2)=(1+2D+...)^2(1-D^2+...)^2(x+1/2)="


"=(1+2D+...)(x+5\/2)=x+9\/2"


"y=c_1e^x+c_2xe^x+c_3cosx+c_4sinx+c_5xcosx+c_6xsinx+x^2e^x\/4+"


"+\frac{1}{32}((x-2)sinx -(x^2+1)cosx+3xcosx+x^2sinx)+x+9/2"


4)

"2x^2yy''+4y^2=x^2(y')^2+2xyy'"


"y=x^2z^2"

then:

"y=x^2z^2"

"y'=2xz^2+2x^2zz'"

"y''=2z^2+4xzz'+4xzz'+2x^2((z')^2+zz'')"


"2x^2x^2z^2(2z^2+4xzz'+4xzz'+2x^2((z')^2+zz''))+4x^4z^4="

"=x^2(2xz^2+2x^2zz')^2+2xx^2z^2(2xz^2+2x^2zz')"


"x^2z^2(z^2+4xzz'+x^2((z')^2+zz''))+x^2z^4="

"=x^2z^4+2x^3z^3z'+x^4z^2(z')^2+xz^2(xz^2+x^2zz')"


"z^2+4xzz'+x^2((z')^2+zz'')+z^2=z^2+2xzz'+x^2(z')^2+z^2+xzz'"

"xzz'+x^2zz''=0"

"xz(z'+xz'')=0"


"z'+xz''=0"

"z'=u"

"u+xu'=0"

"du\/u=-dx\/x"

"lnu=-lnx+lnc_1"

"u=c_1\/x"

"dz=c_1dx\/x"

"z=c_1lnx+c_2"


"y(x)=x^2(c_1lnx+c_2)^2"